OpenStudy (anonymous):

Solve the system of equations: 4x - 3y = -19 2x + y = 13 trying to find the values of x and y

OpenStudy (lgbasallote):

multiply the second equation by 3 4x - 3y = -19 6x + 3y = 39 now add them 4x - 3y = -19 6x + 3y = 39 ========= 10x = 20 solve for x

OpenStudy (anonymous):

OH!